# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 3.3. Random dynamical system
Using a random transformation (known as Doss-Sussman transformation, [29, 59]), we get a random partial differential equation which equivalent to system (2.18). Let us define
@@
$\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}&=-\underbrace{\nabla(\textbf{P}-{p})}_{:=\nabla\widehat{\textbf{P}}}+\nu\Delta\boldsymbol{y}-\big{(}(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla\big{)}(\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})&&\\ &\quad+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))+\chi\boldsymbol{z}+\boldsymbol{f}&&\text{in}\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{y}&=0\quad&&\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}&=\boldsymbol{0}&&\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}(x,0)&=\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0)=:\boldsymbol{y}_{0}\quad&&\text{in }\mathscr{O},\end{aligned}\right.$ (3.18)
@@
[…]
@@
Assume that $\boldsymbol{y}_{0}\in\mathbb{H}$, $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. A function $\boldsymbol{y}$ is called a weak solution of system (3.19) on the time interval $[0,\infty)$, if
@@
    for any $\boldsymbol{\phi}\in\mathbb{V},$
    $\frac{\mathrm{d}\boldsymbol{y}(t)}{\mathrm{d}t},\boldsymbol{\phi}\right>$
@@
Assume that (1.4) is satisfied, $\chi\geq 0$, $\boldsymbol{v}_{0}\in\mathbb{H}$, $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$. Then, there exists a unique weak solution $\boldsymbol{y}(\cdot)$ to system (3.19) in the sense of Definition 3.2 which satisfies the following energy equality:
@@
Let us fix $T>0$. Note that it is enough to prove the result on the interval $[0,T]$. The proof is divided into the following four steps. Step I. Finite-dimensional approximation and weak limits. Let $\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m},\ldots\}$ be a complete orthonormal system in $\mathbb{H}$ belonging to $\mathbb{V}_{s}\subset\mathbb{X}$ (where $s>2$) and let $\mathbb{H}_{m}:=\mathrm{span}\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}\}$. Let $\mathrm{P}_{m}$ denote the orthogonal projection of $\mathbb{X}^{\prime}$ to $\mathbb{H}_{m}$, that is,
$\mathrm{P}_{m}\boldsymbol{x}=\sum_{i=1}^{m}\langle\boldsymbol{x},\boldsymbol{e}_{i}\rangle\boldsymbol{e}_{i}.$ (3.21)
Since every element $\boldsymbol{x}\in\mathbb{H}$ induces a functional $\boldsymbol{x}^{*}\in\mathbb{H}$ by the formula $\langle\boldsymbol{x}^{*},\widetilde{\boldsymbol{x}}\rangle=(\boldsymbol{x},\widetilde{\boldsymbol{x}})$, $\widetilde{\boldsymbol{x}}\in\mathbb{X}$, then $\mathrm{P}_{m}\big{|}_{\mathbb{H}}$, the orthogonal projection of $\mathbb{H}$ onto $\mathbb{H}_{m}$ is given by $\mathrm{P}_{m}\boldsymbol{x}=\sum_{i=1}^{m}\left(\boldsymbol{x},\boldsymbol{e}_{i}\right)\boldsymbol{e}_{i}$. Hence in particular, $\mathrm{P}_{m}$ is the orthogonal projection from $\mathbb{H}$ onto $\mathbb{H}_{m}$.
Let us consider the following system of ODEs on the finite dimensional space $\mathbb{H}_{n}$:
[…]
where
@@
Since $\mathscr{A}_{n}\cdot+\mathscr{B}_{n}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}_{n}(\cdot+\boldsymbol{z})+\beta\mathscr{K}_{n}(\cdot+\boldsymbol{z})$ is locally Lipschitz (see Subsection 2.2), system (3.22) has a unique local solution $\boldsymbol{y}^{m}\in\mathrm{C}([0,T^{*}];\mathbb{H}_{n})$, for some $0<T^{*}<T$. The following uniform estimates show that the time $T^{*}$ can be extended to time $T$. Taking the inner product with $\boldsymbol{y}^{m}(\cdot)$ to the first equation of (3.22) and using (2.4), we obtain
@@
$|\displaystyle\left|\frac{\alpha}{2}\int_{\mathscr{O}}(\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y}^{m})\mathrm{d}x\right|$$\leq\frac{\left|\alpha\right|}{2}\int_{\mathscr{O}}\left|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\right|^{2}\left|\mathrm{A}(\boldsymbol{y}^{m})\right|\mathrm{d}x$
@@
and
$\big{|}\langle\boldsymbol{f},\boldsymbol{y}^{m}\rangle\big{|}$
$=\big{|}\langle\boldsymbol{f}_{1}+\boldsymbol{f}_{2},\boldsymbol{y}^{m}\rangle\big{|}$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\|\boldsymbol{y}^{m}\|_{\mathbb{W}^{1,4}}$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\boldsymbol{y}^{m}\|_{{4}}+\|\nabla\boldsymbol{y}^{m}\|_{{4}})$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\mathrm{A}(\boldsymbol{y}^{m})\|^{\frac{d}{4+d}}_{{4}}\|\boldsymbol{y}^{m}\|_{2}^{\frac{4}{4+d}}+\|\mathrm{A}(\boldsymbol{y}^{m})\|_{{4}})$
$\leq C\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\mathrm{A}(\boldsymbol{y}^{m})\|^{\frac{d}{4+d}}_{{4}}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}^{\frac{4}{4+d}}+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|_{{4}}+\|\mathrm{A}(\boldsymbol{z})\|_{{4}})$
$\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4}\bigg{]}.$ (3.28)
@@
$\leq C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]},$ (3.29)
@@
$\leq\|\boldsymbol{y}^{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]}\mathrm{d}s.$ (3.30)
Hence, an application of Gronwall’s inequality and the fact that $\|\boldsymbol{y}^{m}(0)\|_{2}\leq\|\boldsymbol{y}_{0}\|_{2}$, $\boldsymbol{f}_{1}\in\mathbb{H}^{-1}(\mathscr{O})$, $\boldsymbol{f}_{2}\in\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}\in\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, we have from (3.3) that
@@
For any arbitrary element $\boldsymbol{\phi}\in\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, using Hölder’s inequality and Lemma 2.1, we have from (3.22) that
@@
$\leq\int_{0}^{T}\bigg{[}\nu\left|(\nabla\boldsymbol{y}^{m}(t),\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t))\right|+\left|b(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t),\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\right|+\chi|(\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t))|$
@@
$\qquad+|\langle\boldsymbol{f},\mathrm{P}_{m}\boldsymbol{\phi}(t)\rangle|\bigg{]}\mathrm{d}t+|\alpha|\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))|^{2}|\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t)|\mathrm{d}x\mathrm{d}t$
$\qquad+\beta\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))|^{3}|\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t)|\mathrm{d}x\mathrm{d}t$
$\leq C\bigg{[}\|\boldsymbol{y}^{m}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}+T^{\frac{1}{2}}\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}+T^{\frac{3}{4}}\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{\boldsymbol{z}}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}\bigg{]}$
×$\qquad\times\|\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}$
which implies that the sequences
@@
$\{\frac{\mathrm{d}\boldsymbol{y}^{m}}{\mathrm{d}t}\}_{m\in\mathbb{N}}$ and $\{\mathscr{G}(\boldsymbol{y}^{m})\}_{m\in\mathbb{N}}$ are bounded in $\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})),$ (3.32)
@@
where $\mathscr{G}(\cdot)$ given by (2.5). Taking (3.31)-(3.32) into account and using the Banach-Alaoglu theorem, we obtain the existence of an element $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ with $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ and $\mathscr{G}_{0}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ such that
[…]
along a subsequence (still denoted by the same symbol). Note that $\boldsymbol{f}_{m}\to\mathscr{P}\boldsymbol{f}$ in $\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}_{m}\to\boldsymbol{z}$ in $\mathrm{L}^{2}(0,T;\mathbb{H})$. Therefore, on passing to limit as $m\to\infty$ in (3.22), the limit $\boldsymbol{y}(\cdot)$ satisfies:

### v2

@@ 3.3. Random dynamical system
Using a random transformation (known as Doss-Sussmann transformation, [30, 67]), we get a random partial differential equation which equivalent to system (2.30). Let us define
@@
$\left\{\begin{aligned} \frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}&=-\underbrace{\nabla(\textbf{P}-{p})}_{:=\nabla\widehat{\textbf{P}}}+\nu\Delta\boldsymbol{y}-\big((\boldsymbol{y}+\boldsymbol{z})\cdot\nabla\big)(\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})&&\\ &\quad+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))+\chi\boldsymbol{z}+\boldsymbol{f}&&\text{in }\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{y}&=0\quad&&\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}&=\boldsymbol{0}&&\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}(x,0)&=\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0)=:\boldsymbol{y}_{0}\quad&&\text{in }\mathscr{O},\end{aligned}\right$. (3.18)
@@
$\left\{\begin{aligned} \frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}&=-\nu\mathscr{A}\boldsymbol{y}-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z})-\alpha\mathscr{J}(\boldsymbol{y}+\boldsymbol{z})-\beta\mathscr{K}(\boldsymbol{y}+\boldsymbol{z})+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f},\\ \boldsymbol{y}(0)&=\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0)=:\boldsymbol{y}_{0}.\end{aligned}\right$. (3.19)
@@
Assume that $\boldsymbol{y}_{0}\in\mathbb{H}$, $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$. A function $\boldsymbol{y}$ is called a weak solution of system (3.19) on the time interval $0,\infty)$, if
@@
    for any $\boldsymbol{\phi}\in\mathbb{X},$
    $\left<\frac{\mathrm{d}\boldsymbol{y}(t)}{\mathrm{d}t},\boldsymbol{\phi}\right$>
@@
Assume that (1.4) is satisfied, $\chi\geq 0$, $\boldsymbol{v}_{0}\in\mathbb{H}$, $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$. Then, there exists a unique weak solution $\boldsymbol{y}(\cdot)$ to system (3.19) in the sense of Definition 3.2 which satisfies the following energy equality:
@@
Let us fix $T>0$. Note that it is enough to prove the result on the interval $[0,T]$. The proof is divided into the following four steps.
Step I. Finite-dimensional approximation and weak limits. For each $n\in\mathbb{N}$, we search for approximate solutions of the form
$\boldsymbol{y}^{m}(x,t)=\sum_{k=1}^{m}g^{m}_{k}(t)\boldsymbol{e}_{k}(x),\;\;\boldsymbol{e}_{k}\in\mathbb{H}_{m},$ (3.21)
where $\{\boldsymbol{e}_{k}\}_{k\in\mathbb{N}}$ is defined in (2.26), $\mathbb{H}_{m}$ is defined in Subsection 2.4 and the coefficients $g^{m}_{1},\ldots,g^{m}_{m}$ are solutions of the following $m$ ordinary differential equations:
$\left\{\begin{aligned} \frac{\mathrm{d}}{\mathrm{d}t}\left(\boldsymbol{y}^{m},\boldsymbol{e}_{j}\right)&=-\nu(\mathscr{A}_{m}\boldsymbol{y}^{m}-\mathscr{B}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})-\alpha\mathscr{J}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})-\beta\mathscr{K}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})+\chi\boldsymbol{z}_{m}+\boldsymbol{f}_{m},\boldsymbol{e}_{j}),\\ (\boldsymbol{y}^{m}(0),\boldsymbol{e}_{j})&=(\mathrm{P}_{m}[\boldsymbol{x}-\boldsymbol{z}(\omega)(0)],\boldsymbol{e}_{j})=:({\boldsymbol{y}_{0}}_{m},\boldsymbol{e}_{j}),\end{aligned}\right$. (3.22)
for $j=1,\cdots,m$, where
@@
Since $\mathscr{A}_{m}\cdot+\mathscr{B}_{m}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}_{m}(\cdot+\boldsymbol{z})+\beta\mathscr{K}_{m}(\cdot+\boldsymbol{z})$ is locally Lipschitz (see Subsection 2.3), system (3.22) has a unique local solution $\boldsymbol{y}^{m}\in\mathrm{C}([0,T^{*}_{m}];\mathbb{H}_{m})$, for some $0<T^{*}_{m}\leq T$. The following uniform estimates show that the time $T^{*}_{m}$ can be extended to time $T$. From the first equation of (3.22) and (2.10), we obtain
@@
$\left|\frac{\alpha}{2}\int_{\mathscr{O}}(\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y}^{m})\mathrm{d}x\right|$ $\leq\frac{\left|\alpha\right|}{2}\int_{\mathscr{O}}\left|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\right|^{2}\left|\mathrm{A}(\boldsymbol{y}^{m})\right|\mathrm{d}x$
@@
$\big|\langle\boldsymbol{f},\boldsymbol{y}^{m}\rangle\big|$ $\leq\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}\leq C\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}$
$\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+C\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}.$ (3.28)
@@
$\leq C\bigg[\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg],$ (3.29)
@@
$\leq\|\boldsymbol{y}^{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg[\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}\bigg]\mathrm{d}s.$ (3.30)
Hence, an application of Gronwall’s inequality and the fact that $\|\boldsymbol{y}^{m}(0)\|_{2}\leq\|\boldsymbol{y}_{0}\|_{2}$, $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$, and $\boldsymbol{z}\in\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, we have from (3.3) that
@@
For any arbitrary element $\boldsymbol{\phi}\in\mathrm{L}^{4}(0,T;\mathbb{U})$, using Hölder’s inequality and Lemma 2.6, we have from (3.22) that
@@
$\leq\int_{0}^{T}\bigg[\nu\left|(\nabla\boldsymbol{y}^{m}(t),\nabla\mathrm{P}_{m}\boldsymbol{\phi}(t))\right|+\left|b(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t),\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\right|+\chi|(\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t))|$
@@
$\qquad+|\langle\boldsymbol{f},\mathrm{P}_{m}\boldsymbol{\phi}(t)\rangle|\bigg]\mathrm{d}t+|\alpha|\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(x,t)+\boldsymbol{z}(x,t))|^{2}|\nabla\mathrm{P}_{m}\boldsymbol{\phi}(x,t)|\mathrm{d}x\mathrm{d}t$
$\qquad+\beta\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(x,t)+\boldsymbol{z}(x,t))|^{3}|\nabla\mathrm{P}_{m}\boldsymbol{\phi}(x,t)|\mathrm{d}x\mathrm{d}t$
$\leq\int_{0}^{T}\bigg[\nu\|\boldsymbol{y}^{m}(t)\|_{\mathbb{V}}\|\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{\mathbb{V}}+\|\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t)\|^{2}_{4}\|\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{\mathbb{V}}+\chi\|\boldsymbol{z}(t)\|_{2}\|\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{2}$
$\qquad+\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}\|\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{\mathbb{V}}\bigg]\mathrm{d}t+|\alpha|\int_{0}^{T}\|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\|^{2}_{4}\|\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{\mathbb{V}}\mathrm{d}t$
$\qquad+\beta\int_{0}^{T}\|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\|^{3}_{4}\|\nabla\mathrm{P}_{m}\boldsymbol{\phi}(t)\|_{4}\mathrm{d}t$
$\leq C\bigg[\|\boldsymbol{y}^{m}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}+\|\boldsymbol{y}^{m}+\boldsymbol{z}\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}+\|\boldsymbol{z}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})}$
$\quad+T^{\frac{1}{2}}\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}$
$\quad+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{3}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\|\nabla\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\bigg]$
$\leq C\bigg[\|\boldsymbol{y}^{m}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}+\|\boldsymbol{y}^{m}+\boldsymbol{z}\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}+\|\boldsymbol{z}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})}+T^{\frac{1}{2}}\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\bigg]$
$\quad\times\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{U})}+C\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{3}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\|\mathrm{P}_{m}\boldsymbol{\phi}\|_{\mathrm{L}^{4}(0,T;\mathbb{U})}$
$\leq C\bigg[\|\boldsymbol{y}^{m}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})}+\|\boldsymbol{y}^{m}+\boldsymbol{z}\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}+\|\boldsymbol{z}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})}+T^{\frac{1}{2}}\|\boldsymbol{f}\|_{\mathbb{H}^{-1}}+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{2}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\bigg]$
$\quad\times\|\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{U})}+C\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{3}_{\mathrm{L}^{4}(0,T;\mathbb{L}^{4})}\|\boldsymbol{\phi}\|_{\mathrm{L}^{4}(0,T;\mathbb{U})},$ (3.32)
which implies that the sequence
@@
$\left\{\frac{\mathrm{d}\boldsymbol{y}^{m}}{\mathrm{d}t}\right\}_{m\in\mathbb{N}}\text{ is bounded in }\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{U}^{\prime}).$ (3.33)
@@
Taking (3.31)-(3.33) into account and using the Banach-Alaoglu theorem, we obtain the existence of an element $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ with $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{U}^{\prime})$ such that
$\boldsymbol{y}^{m}\xrightharpoonup{w^{*}}$ $\ \boldsymbol{y}$ $\text{ in }\ \ \ \ \ \mathrm{L}^{\infty}(0,T;\mathbb{H}),$ (3.34)
$\boldsymbol{y}^{m}\xrightharpoonup{w}$ $\ \boldsymbol{y}$ $\text{ in }\ \ \ \ \ \mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$ (3.35)
$\frac{\mathrm{d}\boldsymbol{y}^{m}}{\mathrm{d}t}\xrightharpoonup{w}$ $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}$ $\text{ in }\ \ \ \ \ \mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{U}^{\prime}),$ (3.36)
along a subsequence (still denoted by the same symbol). In addition, we have that the sequence
$\{\mathscr{G}(\boldsymbol{y}^{m})\}_{m\in\mathbb{N}}$ is bounded in $\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})),$ (3.37)
where $\mathscr{G}(\cdot)$ given by (2.11), which implies that there exists $\mathscr{G}_{0}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ such that